A chocolate drink is 16% pure chocolate by volume. If 13 litres of pure milk are added to 22 litres of this drink, the percent of chocolate in the new drink is approximately
Answer by lwsshak3(11628)

A chocolate drink is 16% pure chocolate by volume. If 13 litres of pure milk are added to 22 litres of this drink, the percent of chocolate in the new drink is approximately.
***
let x= percent of chocolate in the new drink.
16%(22)=x%(13+22)
3.52=x(35)
x=3.52/35=0.10
percent of chocolate in the new drink≈10%